Miguel just retired, and has $590,000 to invest. A very safe Certificate of Deposit (CD) account pays 2%, while a riskier bond fund pays 5% in interest. Miguel figures he needs $21,000 a year in interest to live on. How much should he invest in each account to make enough interest while minimizing his risk?
$
at 2%
$
at 5%

Answer by Boreal(15235)

x at 2%=.02x interest a year
590-x at 5%=29.5-.05x a year (will work in thousands)
.02x+29.5-0.05x=21
-0.03x=-8.5
x=283.33
so $283,333.33 at 2%=$5666.67 in interest
and $306,666.67 at 5%=$15,333.33 in interest